OpenStudy (lilai3):

One letter is randomly selected from the word PRIME and one letter is randomly selected from the word MATH. What is the probability that getting both letters selected are vowels? A card is chosen at random from a deck of 52 cards. It is then replaced and a second card is chosen. What is the probability that both letters selected are vowels?

OpenStudy (anonymous):

2/5 * 1/5 = 2/25. 4/52 = 1/13.

OpenStudy (lilai3):

thank you very much and hafa good nite, @scarydoor!

OpenStudy (lilai3):

but waiitttt

OpenStudy (lilai3):

wut about the second one?

OpenStudy (anonymous):

Oh I misread the second one. The chance is 1/13 * 1/13. Since both random events are independent of each other.

OpenStudy (lilai3):

it's 1/169

OpenStudy (lilai3):

so is it 1\169???

OpenStudy (anonymous):

what is 1/13 * 1/13? It's 1/169. In original quesiton, you say what is the chance of choosing a vowel in the card selected. I took that to mean "what is the chance of picking an ace card" since no others have a vowel on their card. So you have 4/52 = 1/13 chance of picking an ace from a pack. Since you choose two cards, from two full decks, they are both independentent of each other. That means that the chance of them both being the ace card is 1/13 * 1/13 = 1/169.

OpenStudy (lilai3):

oh thank you and good night and may da bedbugs not bite u.. teehee :D